Institut for Matematik og Datalogi Odense Universitet March 30, 1995

JFB

DM15 – F95 – Week 24

Lecture, March 27

We finished chapter 12, and sections 13.1-13.3 and 13.6-13.7 of chapter 13.

Lecture, April 3

We will finish chapter 13 in Sudkamp, including the sections skipped last week, and we will cover the first chapter of Mitchell’s monograph.

Lecture, April 10

We will begin on the second chapter of Mitchell’s monograph.

Problems to be discussed in week 16

• Do problem 3 on page 332 of Sudkamp.

• Do problems 6b and 6e on page 333 of Sudkamp.

• Do problem 8 on page 333 of Sudkamp.

• Do problem 12a on page 334 of Sudkamp.

• Do problem 16 on page 335 of Sudkamp.

• Do problem 20 on page 336 of Sudkamp.

• Do exercises 1.2.1, 1.2.2, and 1.2.3 on page 19 of Mitchell.